OpenStudy (anonymous):

Use Gauss's approach not formula to find the sum of 26+27+28+.......+175

OpenStudy (skullpatrol):

Any ideas what "Gauss's approach" was?

OpenStudy (amoodarya):

a=26 + 27 + 28 +.......+175 a=175+174+173 + ...+26 sum of them then you have 2a so divide by 2

OpenStudy (amoodarya):

2a=150*(26+175)